# Particle transport and deposition in wall-sheared thermal turbulence

### v1

@@ 2.4 Simulation settings
We present 2-D simulation results in the cell with $L=2H$ for Rayleigh number in the range of $10^{7}\leq Ra\leq 10^{9}$ and a fixed Prandtl number of $Pr=0.71$ (motivated by the studies of sand grains and pollutants dispersed in the air), while the wall-shear Reynolds number $Re_{w}$ is in the range $0\leq Re_{w}\leq 12000$. To demonstrate the relevance of our results for 3-D wall-sheared thermal convection, we conducted a set of 3-D simulations in a cell with dimensions $L\times H\times W=2\pi h\times 2h\times\pi h$ at $Ra=10^{7}$ and $Pr=0.71$. Here, $h=H/2$ is the half height of the cell. The wall-shear Reynolds number for the 3-D simulations is in the range $0\leq Re_{w}\leq 4000$. Recently, Jie et al. (2022); Motoori et al. (2022); Gao et al. (2023); Zhang et al. (2023) found that this domain size can accurately capture particle behaviors. The properties of air at a reference temperature of 300 K, including density, thermal expansion coefficient, and kinematic viscosity, are listed in table 1. Additionally, we provide the height of the convection cell for a temperature difference of 5 K.
@@
Thermal expansion coefficient ($\beta$) 3.36$\times$10-3 K-1
Kinematic viscosity (${\nu_{f}}$) 1.57$\times$10-5 m2/s
Thermal diffusivity (${\alpha_{f}}$) 2.22$\times$10-5 m2/s
@@
In the Nek5000 solver, the effective grid number is determined by the product of the number of spectral elements and the polynomial order. For the 2-D simulations, the cases with $Ra=10^{7}$ and $10^{8}$ use uniform grids, while the cases with $Ra=10^{9}$ use non-uniform grids. For the 3-D simulations, the $Ra=10^{7}$ cases also use non-uniform grids. Specifically, the mesh spacing in the wall-parallel directions remains uniform for all cases; in the wall-normal direction, we adopt a cosine stretching function $y(\xi)=-\cos[\pi(\xi+1)/2]$ with $\xi\in[-1,1]$ to cluster points (Bernardini et al., 2014). To validate the resolution, grid spacing in the wall-normal direction ${\Delta_{y}}$ and time interval ${\Delta_{t}}$ are compared against the Kolmogorov and Batchelor scales. The Kolmogorov length scale is estimated by the global criterion $\eta_{K}=\left(\nu_{f}^{3}/\langle\varepsilon_{u}\rangle_{V,t}\right)^{1/4}$, the Batchelor length scale is estimated by $\eta_{B}=\eta_{K}Pr^{-1/2}$ (Batchelor, 1959; Silano et al., 2010), and the Kolmogorov time scale is estimated by $\tau_{\eta_{K}}=\sqrt{\nu_{f}/\langle\varepsilon_{u}\rangle_{V,t}}$. Here, $\varepsilon_{u}$ denotes the turbulent kinetic energy dissipation rates, and $\langle\cdots\rangle_{V,t}$ denotes the volume- and time-averaged values. As shown in table 2, the grid spacing in the wall-normal direction meets the criterion $\max({\Delta_{y}}/{\eta_{K}},{\Delta_{y}}/{\eta_{B}})\approx 1$, ensuring sufficient spatial resolution. Additionally, the time intervals satisfy $\max({\Delta_{t}}/\tau_{{\eta_{K}}})\approx 0.007$, guaranteeing adequate temporal resolution.
@@
We have varied both the diameter and the density of the particles in our simulations. Specifically, we explored the parameter space of 5 $\mu$m $\leq{d_{p}}\leq$ 80 $\mu$m and 600 kg/m3 $\leq{\rho_{p}}\leq$ 3000 kg/m3. Initially, the 2-D cell contains $N_{0}=100\times 50=5000$ particles, and the 3-D cell contains $N_{0}=150\times 50\times 75=562500$ particles. Motivated by studies of sand grains and pollutants dispersed in the air ($\rho_{p}/\rho_{f}\gg 1$), we have the buoyancy ratio $\Lambda\approx 1/(\beta\Delta_{T})$ for all the simulation cases. In our simulations, we fixed the temperature difference $\Delta_{T}=5$ K, leading to $\Lambda\approx 59.5$. Then, (14) implies that we have one nondimensional parameter, the Stokes number, to control the particle dynamics. The Stokes number based on the free-fall time scale is defined as $S{t_{f}}={\tau_{p}}/{t_{f}}={\rho_{p}}d_{p}^{2}/\left({18{\mu_{f}}}\right)/\sqrt{H/(g\beta{\Delta_{T}})}$, and the $St_{f}$ characterizes the large-scale effects. On the other hand, the Stokes number based on the Kolmogorov time scale is defined as $S{t_{K}}={\tau_{p}}/{\tau_{{\eta_{K}}}}={\rho_{p}}d_{p}^{2}/\left({18{\mu_{f}}}\right)/\sqrt{\nu_{f}/{{\left\langle{{\varepsilon_{u}}}\right\rangle}_{V,t}}}$, and the $St_{K}$ characterizes the small-scales effects. For clarity in the following discussion, $St$ will denote $St_{K}$ to reflect the particle inertia relative to that of small-scale fluid motion unless otherwise mentioned. We will use particle density ${\rho_{p}}$ and particle diameter ${d_{p}}$ to distinguish between different simulation cases. Dimensional numerical values for various particle quantities of interest in terms of $St_{f}$ and $St_{K}$ are tabulated in the supplementary table. In addition, we estimate the Kolmogorov length scale as ${\eta_{K}}\geq$ 2.0 mm (see table 2), which is around 25 times larger than the largest particle. The largest particle volume fraction of all cases is $\phi\sim\mathcal{O}(10^{-5})$, thus justifying the use of a one-way coupling approach to model the motion of point particles in this work. Even for large $St$, where sedimentation dominates over convective or shear motions, the back reaction of particles on the fluid is generally less significant if $\eta_{K}/d_{p}\gtrsim 10$ (Gore & Crowe, 1989). In appendix A, we further examine the spatial distribution of the time-averaged local Kolmogorov length scale to support that the point-particle model can be safely used. However, we acknowledge that this simplification holds primarily in dilute regimes (Balachandar & Eaton, 2010); in more concentrated suspensions, the back reaction can become important and should be considered.
@@
Table 2: A posteriori check of spatial and temporal resolutions of the simulations. The columns from left to right indicate the following: Rayleigh number $Ra$, wall-shear Reynolds number $Re_{w}$, Richardson number $Ri$, the effective grid number (i.e., the product of spectral element number and polynomial order), the ratio of maximum grid spacing in the wall-normal direction over the Kolmogorov length scale, the ratio of maximum grid spacing in the wall-normal direction over the Batchelor length scale, the ratio of maximum time interval over the Kolmogorov time scale, the Kolmogorov length scale. Note that for the 2-D simulations at $Ra=10^{7}$ with $Re_{w}=2000$, and at $Ra=10^{8}$ with $Re_{w}=8000$, the flow is in the laminar state.
@@ 3.1 Particle transport in the convection cell
In figure 2, we show snapshots of temperature fields at various wall-shear Reynolds numbers $Re_{w}$ from 2-D simulations, and the corresponding video can be viewed in supplementary movie 1. These snapshots are taken when half of the total particles remain suspended in the convection cell. The canonical RB convection (see figures 2a-c) consists of two large-scale rolls rotating in opposite directions. At lower $Re_{w}$ (see figures 2d-f), the horizontal wall shear effects are weak, and the dominant flow patterns are similar to those in the canonical RB convection. As the wall shear increases, the large-scale rolls undergo horizontal expansion, eventually leading to a transition into the zonal flow state (see figures 2g-i). Zonal flows are primarily horizontal flows typically aligned with the rotation axis in rotating systems (Goluskin et al., 2014; Zhang et al., 2020; Lin & Jackson, 2021) or along a specific direction in non-rotating systems (Wang et al., 2020b; Winchester et al., 2021; Jin et al., 2022). These flows exhibit large-scale, coherent structures that can span the entire system and are often associated with banded structures in the flow field. Previous 2-D studies on zonal flow were primarily conducted in convection cells with free-slip boundaries, which do not exert shear stress to slow down the fluid, and the periodicity of the flow system allows for a horizontal mean flow (Goluskin et al., 2014; Wang et al., 2020b; Winchester et al., 2021). In contrast, our work shows that moving boundaries induce a horizontal mean flow, forming the zonal flow. The overall trend in the 2-D flow pattern is similar to that in convection cells with rough shearing walls (Jin et al., 2022).
@@
Figure 2: Typical instantaneous temperature fields (contours) and particle positions (black dots, sizes artificially increased for better flow visualization) from 2-D simulations at (a-c) $Re_{w}$ = 0, (d-f) $Re_{w}$ = 2000, and (g-i) $Re_{w}$ = 6000, with $Ra=10^{8}$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3, with diameters of (a,d,g) ${d_{p}}=5$ $\mu$m, (b,e,h) ${d_{p}}=20$ $\mu$m, and (c,f,i) ${d_{p}}=80$ $\mu$m. Corresponding Stokes numbers are (a) $St=4.86\times 10^{-4}$, (b) $St=7.78\times 10^{-3}$, (c) $St=1.25\times 10^{-1}$, (d) $St=6.64\times 10^{-4}$, (e) $St=1.06\times 10^{-2}$, (f) $St=1.70\times 10^{-1}$, (g) $St=3.93\times 10^{-4}$, (h) $St=6.29\times 10^{-3}$, and (i) $St=1.01\times 10^{-1}$.
@@
Regarding particle movement, the combined effects of particle diameter and particle density are reflected in the dimensionless Stokes number. For clarity of presentation, most results in this paper are shown at a fixed particle density but with varying particle diameters to present small, medium, and large Stokes number cases. Particles with small $St$ (see figures 2a,d,g) are randomly distributed throughout the convection cell. These particles quickly adapt to changes in the surrounding flow, with little to no aggregation. For medium $St$ (see figures 2b,e,h), we observe a greater concentration of particles over upwelling areas, with downwelling regions displaying fewer particles. This counterintuitive observation aligns with our previous findings (Xu et al., 2020) and independent research by Patočka et al. (2020, 2022). The physical interpretation is that for particles circulating within large-scale rolls in the convection cell, when they approach the bottom boundary beneath downwellings, they are likely to be carried by the fast-moving edges of the rolls toward regions beneath upwellings. In these upwelling regions, particles may enter areas with slower flow due to small-scale irregularities caused by new plume generation. These low-velocity regions increase the likelihood of particle sedimentation, leading to an accumulation of particles beneath the upwellings. In sheared convection cells, horizontal shear forces facilitate particle lateral dispersion, which is particularly apparent in figure 2(h). For large $St$ (see figures 2c,f,i), the inertia of the particles is dominant, causing them to settle and become nearly unresponsive to the carrier flow, similar to behavior observed in quiescent fluid environments. As we will demonstrate later, even these large $St$ particles are slightly influenced by the turbulent flow.
@@
In figure 3, we further show snapshots of the velocity component $v$ at the channel center plane and particle positions from 3-D simulations, and the corresponding video can be viewed in supplementary movie 2. In thermal convection, rising fluids are warmer and falling fluids are colder. At $Re_{w}=0$ and $Ra=10^{7}$, the flow structure resembles the canonical RB convection (Stevens et al., 2018). As $Re_{w}$ increases to 2000, the flow enters a transitional regime where the interaction between the imposed shear and buoyancy becomes significant. Large, elongated thermal plumes start to transform into thin, straight, elongated streaks aligned in the streamwise direction. Eventually, streamwise-oriented rolls are formed at $Re_{w}=4000$ (Blass et al., 2020; Pirozzoli et al., 2017). To further elucidate the nature of these large-scale rolls, we present time-averaged streamlines colored by the vertical velocity $v$, as shown in figures 4(a-c). At $Re_{w}=0$, the averaged streamlines show a well-defined roll structure aligned in the spanwise direction. As $Re_{w}$ increases to 4000, the averaged streamlines transition to an organized pattern aligned with the streamwise direction. Additionally, we measure the velocity orientation $\theta$, defined as the angle between the projection of the velocity vector onto the $x-z$ plane and the positive $x$-axis. We calculate the probability density functions (p.d.f.s) of $\theta$ for fluid nodes in the region $0.1H\leq y\leq 0.9H$, and present their time evolution in figures 4(d-f). At $Re_{w}=0$, the p.d.f.s indicate high probability values consistently around $0^{\circ}$ and $180^{\circ}$, suggesting that the roll axes are predominantly aligned in the spanwise direction, without exhibiting complex motions as seen in an RB cell with a larger length-to-width aspect ratio (Vogt et al., 2018; Li et al., 2022; Teimurazov et al., 2023). At $Re_{w}=4000$, the p.d.f.s shift to show high probability values around ±$\pm 90^{\circ}$, indicating that the roll axes are consistently aligned in the streamwise direction.
@@
Figure 3: Typical instantaneous slice of velocity component $v$ at the channel center plane (contours) and particle positions (black dots, sizes artificially increased for better flow visualization, here 0.8 % of suspended particles are shown for clarity) from 3-D simulations at (a-c) $Re_{w}$ = 0, (d-f) $Re_{w}$ = 2000, and (g-i) $Re_{w}$ = 4000, with $Ra=10^{7}$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3, with diameters of (a,d,g) ${d_{p}}=5$ $\mu$m, (b,e,h) ${d_{p}}=20$ $\mu$m, and (c,f,i) ${d_{p}}=80$ $\mu$m. Corresponding Stokes numbers are (a) $St=7.05\times 10^{-4}$, (b) $St=1.13\times 10^{-2}$, (c) $St=1.80\times 10^{-1}$, (d) $St=7.70\times 10^{-4}$, (e) $St=1.23\times 10^{-2}$, (f) $St=1.97\times 10^{-1}$, (g) $St=8.83\times 10^{-4}$, (h) $St=1.41\times 10^{-2}$, and (i) $St=2.26\times 10^{-1}$.
@@
Figure 4: (a-c) Time-averaged streamlines colored by the vertical velocity $v$, (d-f) time evolution of the probability density functions (p.d.f.s) of the instantaneous velocity orientation $\theta$ at (a,d) $Re_{w}$ = 0, (b,e) $Re_{w}$ = 2000, and (c,f) $Re_{w}$ = 4000, with $Ra=10^{7}$.
@@
We then examine the global response parameters of the Nusselt number ($Nu$) and Reynolds number ($Re$) on the control parameter $Re_{w}$ in the 2-D flow system. Here, the heat transfer efficiency is calculated as $Nu=1+\sqrt{PrRa}{\langle v^{*}T^{*}\rangle_{V,t}}$, and the global flow strength of the convection is calculated as $Re=\sqrt{{{\langle\|\mathbf{u}\|^{2}\rangle}_{V,t}}}H/\nu_{f}$. Moreover, we consider the lateral and vertical components of the $Re$, respectively, as $Re_{x}=\sqrt{{{\langle{u^{2}}\rangle}_{V,t}}}H/\nu_{f}$, $Re_{y}=\sqrt{{{\langle{v^{2}}\rangle}_{V,t}}}H/\nu_{f}$. Both $Nu$ and $Re$ exhibit sharp changes and abrupt transitions at $Re_{w}\simeq 1500$ with $Ra=10^{7}$ (see figures 5a and 5d), at $Re_{w}\simeq 5000$ with $Ra=10^{8}$ (see figures 5b and 5e), suggesting a threshold for the competition between wall shear and buoyancy that leads to changes in the flow state from LSCs to zonal flows in 2-D simulations. The heat transport behavior in wall-sheared thermal turbulence is governed by the interaction between the buoyancy-driven LSC and the shear introduced by the moving walls. At higher wall shear, strong wall instabilities disrupt the large-scale rolls, which are crucial for carrying thermal energy away from the wall and efficiently transporting heat. When these large-scale structures are disrupted, small-scale turbulent motions become more prevalent, and the convective motions near the wall increase. However, these motions become more localized and less effective in transporting heat across larger vertical distances, thus decreasing the overall efficiency of heat transport. Regarding the velocity fields, LSCs exhibit both significant vertical and horizontal velocity components, while zonal flows in 2-D show strong horizontal velocities with a less vertical component. At a fixed Rayleigh number, when the wall shear becomes sufficiently strong, the flow transitions to laminar Couette flow, causing the vertical flow velocity component to become nearly zero. For a Rayleigh number of $10^{7}$, this transition occurs approximately at $Re_{w}\simeq 2000$ (see figure 5d); for a Rayleigh number of $10^{8}$, the transition happens at around $Re_{w}\simeq 8000$ (see figure 5e). At an even higher Rayleigh number of $10^{9}$, based on the results of Jin et al. (2022) and Xu et al. (2023), we speculate that the transitions to laminar Couette flow occurs at a much larger $Re_{w}$, beyond the parameter range explored in the present work.

### v2

@@ 2.4 Simulation settings
We present 2-D simulation results in the cell with $L=2H$ for Rayleigh numbers in the range of $10^{7}\leq Ra\leq 10^{9}$ and a fixed Prandtl number of $Pr=0.71$ (motivated by the studies of sand grains and pollutants dispersed in the air), while the wall-shear Reynolds number $Re_{w}$ is in the range $0\leq Re_{w}\leq 12000$. To demonstrate the relevance of our results for 3-D wall-sheared thermal convection, we conducted a set of 3-D simulations in a cell with dimensions $L\times H\times W=2\pi h\times 2h\times\pi h$ at $Ra=10^{7}$ and $Pr=0.71$. Here, $h=H/2$ is the half-height of the cell. The wall-shear Reynolds number for the 3-D simulations is in the range $0\leq Re_{w}\leq 4000$. Recently, Jie et al. (2022), Motoori et al. (2022), Gao et al. (2023) and Zhang et al. (2023) found that this domain size can accurately capture particle behaviours. The properties of air at a reference temperature of 300 K, including density, thermal expansion coefficient and kinematic viscosity, are listed in table 1. Additionally, we provide the height of the convection cell for a temperature difference of 5 K.
@@
Thermal expansion coefficient ($\beta$) $3.36\times 10^{-3}$ K-1
Kinematic viscosity (${\nu_{f}}$) $1.57\times 10^{-5}$ m2/s
Thermal diffusivity (${\alpha_{f}}$) $2.22\times 10^{-5}$ m2/s
@@
In the Nek5000 solver, the effective grid number is determined by the product of the number of spectral elements and the polynomial order. For the 2-D simulations, the cases with $Ra=10^{7}$ and $10^{8}$ use uniform grids, while the cases with $Ra=10^{9}$ use non-uniform grids. For the 3-D simulations, the $Ra=10^{7}$ cases also use non-uniform grids. Specifically, the mesh spacing in the wall-parallel directions remains uniform for all cases; in the wall-normal direction, we adopt a cosine stretching function $y(\xi)=-\cos[\pi(\xi+1)/2]$ with $\xi\in[-1,1]$ to cluster points (Bernardini et al., 2014). To validate the resolution, grid spacing in the wall-normal direction ${\Delta_{y}}$ and time interval ${\Delta_{t}}$ are compared against the Kolmogorov and Batchelor scales. The Kolmogorov length scale is estimated by the global criterion $\eta_{K}=\left(\nu_{f}^{3}/\langle\varepsilon_{u}\rangle_{V,t}\right)^{1/4}$, the Batchelor length scale is estimated by $\eta_{B}=\eta_{K}Pr^{-1/2}$ (Batchelor, 1959; Silano et al., 2010) and the Kolmogorov time scale is estimated by $\tau_{\eta_{K}}=\sqrt{\nu_{f}/\langle\varepsilon_{u}\rangle_{V,t}}$. Here, $\varepsilon_{u}$ denotes the turbulent kinetic energy dissipation rates, and $\langle\cdots\rangle_{V,t}$ denotes the volume- and time-averaged values. As shown in table 2, the grid spacing in the wall-normal direction meets the criterion $\max({\Delta_{y}}/{\eta_{K}},{\Delta_{y}}/{\eta_{B}})\approx 1$, ensuring sufficient spatial resolution. Additionally, the time intervals satisfy $\max({\Delta_{t}}/\tau_{{\eta_{K}}})\approx 0.007$, guaranteeing adequate temporal resolution.
@@
We have varied both the diameter and the density of the particles in our simulations. Specifically, we explored the parameter space of 5 $\mu$m $\leq{d_{p}}\leq$ 80 $\mu$m and 600 kg/m3 $\leq{\rho_{p}}\leq$ 3000 kg/m3. Initially, the 2-D cell contains $N_{0}=100\times 50=5000$ particles, and the 3-D cell contains $N_{0}=150\times 50\times 75=562500$ particles. Motivated by studies of sand grains and pollutants dispersed in the air ($\rho_{p}/\rho_{f}\gg 1$), we have the buoyancy ratio $\Lambda\approx 1/(\beta\Delta_{T})$ for all the simulation cases. In our simulations, we fixed the temperature difference $\Delta_{T}=5$ K, leading to $\Lambda\approx 59.5$. Then, (14) implies that we have one non-dimensional parameter, the Stokes number, to control the particle dynamics. The Stokes number based on the free-fall time scale is defined as $S{t_{f}}={\tau_{p}}/{t_{f}}={\rho_{p}}d_{p}^{2}/\left({18{\mu_{f}}}\right)/\sqrt{H/(g\beta{\Delta_{T}})}$, and $St_{f}$ characterizes the large-scale effects. On the other hand, the Stokes number based on the Kolmogorov time scale is defined as $S{t_{K}}={\tau_{p}}/{\tau_{{\eta_{K}}}}={\rho_{p}}d_{p}^{2}/\left({18{\mu_{f}}}\right)/\sqrt{\nu_{f}/{{\left\langle{{\varepsilon_{u}}}\right\rangle}_{V,t}}}$, and $St_{K}$ characterizes the small-scale effects. For clarity in the following discussion, $St$ will denote $St_{K}$ to reflect the particle inertia relative to that of small-scale fluid motion unless otherwise mentioned. We will use particle density ${\rho_{p}}$ and particle diameter ${d_{p}}$ to distinguish between different simulation cases. Dimensional numerical values for various particle quantities of interest in terms of $St_{f}$ and $St_{K}$ are tabulated in the supplementary table available at https://doi.org/10.1017/jfm.2024.936. In addition, we estimate the Kolmogorov length scale as ${\eta_{K}}\geq$ 2.0 mm (see table 2), which is around 25 times larger than the largest particle. The largest particle volume fraction of all cases is $\phi\sim\mathcal{O}(10^{-5})$, thus justifying the use of a one-way coupling approach to model the motion of point particles in this work. Even for large $St$, where sedimentation dominates over convective or shear motions, the back reaction of particles on the fluid is generally less significant if $\eta_{K}/d_{p}\gtrsim 10$ (Gore & Crowe, 1989). In Appendix A, we further examine the spatial distribution of the time-averaged local Kolmogorov length scale to support that the point-particle model can be safely used. However, we acknowledge that this simplification holds primarily in dilute regimes (Balachandar & Eaton, 2010); in more concentrated suspensions, the back reaction can become important and should be considered.
@@
Table 2: An a posteriori check of spatial and temporal resolutions of the simulations. The columns from left to right indicate the following: Rayleigh number $Ra$, wall-shear Reynolds number $Re_{w}$, Richardson number $Ri$, the effective grid number (i.e. the product of spectral element number and polynomial order), the ratio of maximum grid spacing in the wall-normal direction over the Kolmogorov length scale, the ratio of maximum grid spacing in the wall-normal direction over the Batchelor length scale, the ratio of maximum time interval over the Kolmogorov time scale, the Kolmogorov length scale. Note that for the 2-D simulations at $Ra=10^{7}$ with $Re_{w}=2000$, and at $Ra=10^{8}$ with $Re_{w}=8000$, the flow is in the laminar state.
@@ 3.1 Particle transport in the convection cell
In figure 2, we show snapshots of temperature fields at various wall-shear Reynolds numbers $Re_{w}$ from 2-D simulations, and the corresponding movie can be viewed in supplementary movie 1. These snapshots are taken when half of the total particles remain suspended in the convection cell. The canonical RB convection (see figures 2a-c) consists of two large-scale rolls rotating in opposite directions. At lower $Re_{w}$ (see figures 2d-f), the horizontal wall-shear effects are weak, and the dominant flow patterns are similar to those in the canonical RB convection. As the wall shear increases, the large-scale rolls undergo horizontal expansion, eventually leading to a transition into the zonal flow state (see figures 2g-i). Zonal flows are primarily horizontal flows typically aligned with the rotation axis in rotating systems (Goluskin et al., 2014; Zhang et al., 2020; Lin & Jackson, 2021) or along a specific direction in non-rotating systems (Wang et al., 2020b; Winchester et al., 2021; Jin et al., 2022). These flows exhibit large-scale, coherent structures that can span the entire system and are often associated with banded structures in the flow field. Previous 2-D studies on zonal flow were primarily conducted in convection cells with free-slip boundaries, which do not exert shear stress to slow down the fluid, and the periodicity of the flow system allows for a horizontal mean flow (Goluskin et al., 2014; Wang et al., 2020b; Winchester et al., 2021). In contrast, our work shows that moving boundaries induce a horizontal mean flow, forming the zonal flow. The overall trend in the 2-D flow pattern is similar to that in convection cells with rough shearing walls (Jin et al., 2022).
@@
Figure 2: Typical instantaneous temperature fields (contours) and particle positions (black dots, sizes artificially increased for better flow visualization) from 2-D simulations at (a-c) $Re_{w}$ = 0, (d-f) $Re_{w}$ = 2000 and (g-i) $Re_{w}$ = 6000, with $Ra=10^{8}$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3, with diameters of (a,d,g) ${d_{p}}=5$ $\mu$m, (b,e,h) ${d_{p}}=20$ $\mu$m and (c,f,i) ${d_{p}}=80$ $\mu$m. Corresponding Stokes numbers are (a) $St=4.86\times 10^{-4}$, (b) $St=7.78\times 10^{-3}$, (c) $St=1.25\times 10^{-1}$, (d) $St=6.64\times 10^{-4}$, (e) $St=1.06\times 10^{-2}$, (f) $St=1.70\times 10^{-1}$, (g) $St=3.93\times 10^{-4}$, (h) $St=6.29\times 10^{-3}$ and (i) $St=1.01\times 10^{-1}$.
@@
Regarding particle movement, the combined effects of particle diameter and particle density are reflected in the dimensionless Stokes number. For clarity of presentation, most results in this paper are shown at a fixed particle density but with varying particle diameters to present small, medium and large Stokes number cases. Particles with small $St$ (see figures 2a,d,g) are randomly distributed throughout the convection cell. These particles quickly adapt to changes in the surrounding flow, with little to no aggregation. For medium $St$ (see figures 2b,e,h), we observe a greater concentration of particles over upwelling areas, with downwelling regions displaying fewer particles. This counterintuitive observation aligns with our previous findings (Xu et al., 2020) and independent research by Patočka et al. (2020, 2022). The physical interpretation is that, for particles circulating within large-scale rolls in the convection cell, when they approach the bottom boundary beneath downwellings, they are likely to be carried by the fast-moving edges of the rolls toward regions beneath upwellings. In these upwelling regions, particles may enter areas with slower flow due to small-scale irregularities caused by new plume generation. These low-velocity regions increase the likelihood of particle sedimentation, leading to an accumulation of particles beneath the upwellings. In sheared convection cells, horizontal shear forces facilitate particle lateral dispersion, which is particularly apparent in figure 2(h). For large $St$ (see figures 2c,f,i), the inertia of the particles is dominant, causing them to settle and become nearly unresponsive to the carrier flow, similar to behaviour observed in quiescent fluid environments. As we will demonstrate later, even these large $St$ particles are slightly influenced by the turbulent flow.
@@
In figure 3, we further show snapshots of the velocity component $v$ at the channel centre plane and particle positions from 3-D simulations, and the corresponding movie can be viewed in supplementary movie 2. In thermal convection, rising fluids are warmer and falling fluids are colder. At $Re_{w}=0$ and $Ra=10^{7}$, the flow structure resembles the canonical RB convection (Stevens et al., 2018). As $Re_{w}$ increases to 2000, the flow enters a transitional regime where the interaction between the imposed shear and buoyancy becomes significant. Large, elongated thermal plumes start to transform into thin, straight, elongated streaks aligned in the streamwise direction. Eventually, streamwise-oriented rolls are formed at $Re_{w}=4000$ (Blass et al., 2020; Pirozzoli et al., 2017). To further elucidate the nature of these large-scale rolls, we present time-averaged streamlines coloured by the vertical velocity $v$, as shown in figures 4(a-c). At $Re_{w}=0$, the averaged streamlines show a well-defined roll structure aligned in the spanwise direction. As $Re_{w}$ increases to 4000, the averaged streamlines transition to an organized pattern aligned with the streamwise direction. Additionally, we measure the velocity orientation $\theta$, defined as the angle between the projection of the velocity vector onto the $x-z$ plane and the positive $x$-axis. We calculate the probability density functions (p.d.f.s) of $\theta$ for fluid nodes in the region $0.1H\leq y\leq 0.9H$, and present their time evolution in figures 4(d-f). At $Re_{w}=0$, the p.d.f.s indicate high probability values consistently around $0^{\circ}$ and $180^{\circ}$, suggesting that the roll axes are predominantly aligned in the spanwise direction, without exhibiting complex motions as seen in an RB cell with a larger length-to-width aspect ratio (Vogt et al., 2018; Li et al., 2022; Teimurazov et al., 2023). At $Re_{w}=4000$, the p.d.f.s shift to show high probability values around ±$\pm 90^{\circ}$, indicating that the roll axes are consistently aligned in the streamwise direction.
@@
Figure 3: Typical instantaneous slice of velocity component $v$ at the channel centre plane (contours) and particle positions (black dots, sizes artificially increased for better flow visualization, here 0.8 % of suspended particles are shown for clarity) from 3-D simulations at (a-c) $Re_{w}$ = 0, (d-f) $Re_{w}$ = 2000 and (g-i) $Re_{w}$ = 4000, with $Ra=10^{7}$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3, with diameters of (a,d,g) ${d_{p}}=5$ $\mu$m, (b,e,h) ${d_{p}}=20$ $\mu$m and (c,f,i) ${d_{p}}=80$ $\mu$m. Corresponding Stokes numbers are (a) $St=7.05\times 10^{-4}$, (b) $St=1.13\times 10^{-2}$, (c) $St=1.80\times 10^{-1}$, (d) $St=7.70\times 10^{-4}$, (e) $St=1.23\times 10^{-2}$, (f) $St=1.97\times 10^{-1}$, (g) $St=8.83\times 10^{-4}$, (h) $St=1.41\times 10^{-2}$ and (i) $St=2.26\times 10^{-1}$.
@@
Figure 4: (a-c) Time-averaged streamlines coloured by the vertical velocity $v$, (d-f) time evolution of the p.d.f.s of the instantaneous velocity orientation $\theta$ at (a,d) $Re_{w}$ = 0, (b,e) $Re_{w}$ = 2000 and (c,f) $Re_{w}$ = 4000, with $Ra=10^{7}$.
@@
We then examine the global response parameters of the Nusselt number ($Nu$) and Reynolds number ($Re$) on the control parameter $Re_{w}$ in the 2-D flow system. Here, the heat transfer efficiency is calculated as $Nu=1+\sqrt{PrRa}{\langle v^{*}T^{*}\rangle_{V,t}}$, and the global flow strength of the convection is calculated as $Re=\sqrt{{{\langle\|\mathbf{u}\|^{2}\rangle}_{V,t}}}H/\nu_{f}$. Moreover, we consider the lateral and vertical components of $Re$, respectively, as $Re_{x}=\sqrt{{{\langle{u^{2}}\rangle}_{V,t}}}H/\nu_{f}$, $Re_{y}=\sqrt{{{\langle{v^{2}}\rangle}_{V,t}}}H/\nu_{f}$. Both $Nu$ and $Re$ exhibit sharp changes and abrupt transitions at $Re_{w}\simeq 1500$ with $Ra=10^{7}$ (see figures 5a,d), at $Re_{w}\simeq 5000$ with $Ra=10^{8}$ (see figures 5b,e), suggesting a threshold for the competition between wall shear and buoyancy that leads to changes in the flow state from LSCs to zonal flows in 2-D simulations. The heat transport behaviour in wall-sheared thermal turbulence is governed by the interaction between the buoyancy-driven LSC and the shear introduced by the moving walls. At higher wall shear, strong wall instabilities disrupt the large-scale rolls, which are crucial for carrying thermal energy away from the wall and efficiently transporting heat. When these large-scale structures are disrupted, small-scale turbulent motions become more prevalent, and the convective motions near the wall increase. However, these motions become more localized and less effective in transporting heat across larger vertical distances, thus decreasing the overall efficiency of heat transport. Regarding the velocity fields, LSCs exhibit both significant vertical and horizontal velocity components, while zonal flows in two dimensions show strong horizontal velocities with a lower vertical component. At a fixed Rayleigh number, when the wall shear becomes sufficiently strong, the flow transitions to laminar Couette flow, causing the vertical flow velocity component to become nearly zero. For a Rayleigh number of $10^{7}$, this transition occurs approximately at $Re_{w}\simeq 2000$ (see figure 5d); for a Rayleigh number of $10^{8}$, the transition happens at around $Re_{w}\simeq 8000$ (see figure 5e). At an even higher Rayleigh number of $10^{9}$, based on the results of Jin et al. (2022) and Xu et al. (2023), we speculate that the transitions to laminar Couette flow occurs at a much larger $Re_{w}$, beyond the parameter range explored in the present work.